OpenStudy (anonymous):

Simplify the following expression: (12x7)/(6x2). A. 2x5 B. 2x9 C. 6x5 D. 6x9

OpenStudy (jdoe0001):

\(\bf\large \cfrac{12x^7}{6x^2}\quad ?\)

OpenStudy (anonymous):

i would imagine thats how its wrote in the assignment

OpenStudy (anonymous):

i think its 6x9

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \cfrac{12x^7}{6x^2}\implies\cfrac{2\times 2\times 3xxxxxxx}{2\times 3 xx}\) see any like-terms to cancel out?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\bf \cfrac{12x^7}{6x^2}\implies\cfrac{2\times \cancel{2}\times \cancel{3}\cancel{xx}xxxxx}{\cancel{2}\times \cancel{3} \cancel{xx}}\) so, what's left?

OpenStudy (anonymous):

2x5

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw